# Distributed Deep Learning with RIS Grouping for Accurate Cascaded Channel Estimation

### v1

@@ I Introduction
Reconfigurable intelligent surface (RIS) panels have emerged as a transformative technology for beyond-fifth-generation (B5G) and sixth-generation (6G) wireless systems. These surfaces consist of low-cost, nearly passive metasurfaces that can programmatically control the wireless propagation environment [1, 4]. RIS is especially beneficial in millimeter-wave (mmWave) and terahertz (THz) frequency bands, where signal blockage and severe path loss are prevalent. By intelligently redirecting reflections, RIS can extend coverage and boost spectral efficiency without requiring active radio-frequency (RF) chains on the surface [2, 3, 5]. A primary challenge in harnessing these benefits is acquiring channel state information (CSI) for the cascaded link involving the base station (BS), RIS, and user. Due to the passive design of RIS elements, the BS-RIS and RIS-user channels cannot be estimated independently; instead, only their cascaded product is observable at the receiver. In frequency-division duplexing (FDD) systems, where uplink-downlink reciprocity does not hold, downlink pilots must activate the RIS, and the user measures the combined reflected signals. For an RIS with $N$ elements and a BS with $M$ antennas, the cascaded channel comprises $NM$ complex parameters. Traditional least-squares (LS) estimation demands at least $Q\geq NM$ pilot symbols (or $Q\geq N^{\prime}M$ with grouping, where $N^{\prime}<N$), which becomes impractical for large $N$ due to excessive overhead [6, 7, 8, 9, 10, 19]. Advanced estimators exploit channel structures, such as angular sparsity and spatial correlations, to reduce pilot requirements. Deep neural networks can map low-dimensional pilot observations to the full high-dimensional cascaded channel, often integrating compressive sensing or denoising techniques [11, 12]. Nevertheless, models trained in isolated user or scenario-specific settings frequently exhibit poor generalization to varying user locations, blockage conditions, or angular spreads. This is exacerbated by non-independent and identically distributed (non-IID) data across a cell, leading to degraded accuracy and reliability [11, 12, 14]. Distributed machine learning (DML), including federated learning (FL), tackles non-IID data by aggregating model updates from distributed devices without centralizing raw data, as in the Federated Averaging (FedAvg) algorithm [17]. For heterogeneous environments, specialization enhances performance: sparsely-gated mixture-of-experts (MoE) models use a learned gate to route inputs to a subset of experts, balancing accuracy and efficiency [18]. In RIS CSI estimation, a gate based on coarse scenario indicators (e.g., elevation sectors) can select region-specific experts, avoiding generic feature extraction. Recent developments, such as transformer-based CSI prediction [15] and comprehensive surveys [14], emphasize the need for estimators that (i) function with few pilots ($Q$), (ii) generalize across regions, and (iii) maintain low online computational cost at user equipment (UE). In this paper, we present a DML framework for downlink cascaded channel estimation under limited pilot regimes. Our approach integrates three key components: (i) RIS grouping to compress the effective cascaded channel from $NM$ to $N^{\prime}M$ parameters (with $N^{\prime}=N/g$ for group size $g$), thereby reducing pilot and computational demands; (ii) a region-gated MoE architecture that classifies the propagation region first and then activates only the relevant expert for feature extraction from the pilot data; and (iii) synchronous FedAvg involving the BS and users to train a shared estimator on diverse, non-IID data without exchanging raw samples [17]. The gate employs sparse routing inspired by [18], ensuring that only one expert is active during inference, maintaining UE runtime comparable to a single compact convolutional neural network (CNN).
@@ I-A Related Work
To handle heterogeneity across users and regions without aggregating raw data, DML aggregates model parameters from devices using algorithms like FedAvg [17]. In RIS contexts, [13] proposed a DML scheme for downlink channel estimation, demonstrating enhanced generalization over per-user training. Compared to [13], our work adds (i) RIS grouping to shrink the parameter space from $NM$ to $N^{\prime}M$, cutting pilot and compute costs; and (ii) a region-gated design that engages only one expert per inference, reducing on-device complexity while preserving accuracy.
@@ I-B Contributions
  • A FedAvg-based DML protocol involving users and the BS to train the shared model and experts on local non-IID data, enhancing generalization across regions [17, 13].
@@ II-D RIS Grouping Model
To reduce pilot overhead and computational complexity, we incorporate RIS element grouping via a fixed selection/averaging matrix $\mathbf{S}\in\{0,1\}^{N^{\prime}\times N}$ that maps $N$ physical elements to $N^{\prime}$ control units ($N^{\prime}=N/g$ for cluster size $g$). This grouping aggregates nearby elements into clusters, preserving total reflection energy while reducing the degrees of freedom. The RIS phase vectors satisfy
$\boldsymbol{\phi}=\mathbf{S}^{\mathrm{T}}\bar{\boldsymbol{\phi}},\qquad\bar{\boldsymbol{\phi}}\in\mathbb{C}^{N^{\prime}\times 1},\ \ |\bar{\phi}_{i}|=1,$ (10)
@@ II-E Problem Formulation
$\underbrace{\mathbf{y}_{r,k}^{\mathrm{p}}}_{\in\mathbb{C}^{Q}}=\underbrace{\boldsymbol{\Psi}_{r,k}}_{\in\mathbb{C}^{Q\times NM}}\underbrace{\mathbf{h}_{r,k}}_{\in\mathbb{C}^{NM}}+\ \mathbf{n}_{r,k},\quad\text{or}\quad\mathbf{y}_{r,k}^{\mathrm{p}}=\bar{\boldsymbol{\Psi}}_{r,k}\bar{\mathbf{h}}_{r,k}+\mathbf{n}_{r,k}.$ (13)
@@
$\hat{\mathbf{h}}_{r,k}=g_{\boldsymbol{\vartheta}}\big(\mathbf{y}_{r,k}^{\mathrm{p}};\boldsymbol{\Psi}_{r,k}\big),\quad\text{or}\quad\hat{\bar{\mathbf{h}}}_{r,k}=g_{\boldsymbol{\vartheta}}\big(\mathbf{y}_{r,k}^{\mathrm{p}};\bar{\boldsymbol{\Psi}}_{r,k}\big).$ (14)
@@
$\mathrm{NMSE}\triangleq\mathbb{E}\!\left[\frac{\|\hat{\mathbf{h}}_{r,k}-\mathbf{h}_{r,k}\|_{2}^{2}}{\|\mathbf{h}_{r,k}\|_{2}^{2}}\right],\;\text{or}\;\mathbb{E}\!\left[\frac{\|\hat{\bar{\mathbf{h}}}_{r,k}-\bar{\mathbf{h}}_{r,k}\|_{2}^{2}}{\|\bar{\mathbf{h}}_{r,k}\|_{2}^{2}}\right].$ (15)
@@
$\hat{\mathbf{h}}_{r,k}^{\mathrm{LS}}=\big(\boldsymbol{\Psi}_{r,k}^{\mathrm{H}}\boldsymbol{\Psi}_{r,k}\big)^{-1}\boldsymbol{\Psi}_{r,k}^{\mathrm{H}}\mathbf{y}_{r,k}^{\mathrm{p}},\;\hat{\bar{\mathbf{h}}}_{r,k}^{\mathrm{LS}}=\big(\bar{\boldsymbol{\Psi}}_{r,k}^{\mathrm{H}}\bar{\boldsymbol{\Psi}}_{r,k}\big)^{-1}\bar{\boldsymbol{\Psi}}_{r,k}^{\mathrm{H}}\mathbf{y}_{r,k}^{\mathrm{p}}.$ (18)
@@
MMSE requires a channel covariance model; we adopt a data-driven covariance estimated from samples in the baselines section (cf. Sec. IV).
@@ III-B3 Feature Mapper
$\hat{\mathbf{h}}_{r,k}=M_{\boldsymbol{\psi}}(\mathbf{Z}_{r,k})\quad\text{or}\quad\hat{\bar{\mathbf{h}}}_{r,k}=M_{\boldsymbol{\psi}}(\mathbf{Z}_{r,k}),$ (22)
@@ III-D Distributed Training Protocol
The DML training employs a synchronous federated learning protocol, with the BS serving as the central parameter server and the UEs as clients. This setup leverages the Federated Averaging (FedAvg) algorithm [17] to aggregate local updates, effectively handling the non-IID nature of channel data across regions. Let $\boldsymbol{\vartheta}=(\{\boldsymbol{\theta}_{r}\}_{r=1}^{R},\boldsymbol{\psi})$ denote the trainable parameters of the experts and mapper. In each communication round, the BS broadcasts the current model, clients perform local gradient computations on their private datasets, and the BS aggregates these updates using data-weighted averaging to update the global model.
@@
12:  Aggregate (FedAvg) at BS with weights $p_{r,k}=|\mathcal{D}_{r,k}|/\sum_{r^{\prime},k^{\prime}}|\mathcal{D}_{r^{\prime},k^{\prime}}|$:
@@ III-F Computational Complexity
$\mathcal{C}_{\mathrm{map}}=\mathcal{O}\big(E_{1}E_{2}\cdot 32\cdot(2NM)\big)\quad\text{or}\quad\mathcal{O}\big(E_{1}E_{2}\cdot 32\cdot(2N^{\prime}M)\big),$ (27)
@@ A. Arrays, Pilots, and SNR
RIS pilots $\boldsymbol{\Theta}$ i.i.d. $\{\pm 1\}$ (unit modulus)22footnotemark: 2
@@
  • Unit-modulus pilots satisfy $|\phi_{q,n}|=1$. If desired, use $\tilde{\mathbf{y}}^{\mathrm{p}}=\mathbf{y}^{\mathrm{p}}/\sqrt{Q}$ for energy normalization without altering (8).
@@ IV-C Ablations and Discussion
Distributed training (DML) vs. centralized/single-user training: FedAvg-based DML outperforms single-region or per-user training, with gains most evident at low SNR where regional differences in noise and angles are pronounced. This collaborative approach leverages non-IID data without privacy risks.
@@ Acknowledgment
The authors would like to thank the anonymous reviewers for their valuable comments and suggestions that helped improve the quality of this paper.
Biographies
Fazal Muhammad received the B.Sc. and M.S. degrees in Electrical Engineering from the University of Engineering and Technology, Peshawar, Pakistan, in 2004 and 2007, respectively, and the Ph.D. degree from the Ghulam Ishaq Khan Institute of Engineering Science and Technology, Topi, Pakistan, in 2017. He is currently working as an Assistant Professor at University of Engineering Technology, Mardan. His research is focused on the modeling and analysis of heterogeneous cellular networks. His other research interests include interference channel modeling, optical networks, and antennas modeling and design for 5G and beyond, blockchain technology and digital twin Technology.
Syed Luqman Shah received his B.Sc. degree in Electrical Engineering from the University of Engineering and Technology, Peshawar, where he was awarded the University Gold Medal (2017-2021). He then earned his M.S. degree in Electrical Engineering from the Ghulam Ishaq Khan Institute of Engineering Sciences and Technology, Pakistan, with distinction (2021-2023). After completing his M.S., he worked as a research assistant at FAST National University of Computer and Emerging Sciences, Islamabad Campus (2023-2024). In 2024, he joined the 6G Flagship as a doctoral researcher. His research interests include statistical learning for critical indoor wireless applications, digital twins for wireless networks, wireless channel modeling, and the application of machine learning in wireless communications.
Salman Khan received his Bachelor’s degree in Electrical Engineering from the University of Engineering and Technology (UET), Mardan, Pakistan, in 2025. His academic interests include advanced wireless communication systems, deep learning, and artificial intelligence. He recently completed a research project on deep learning-based channel estimation techniques for 6G networks, focusing on the integration of neural networks with Massive MIMO and Reconfigurable Intelligent Surfaces (RIS). His research interests include the application of machine learning to wireless communication, optimization of 6G and beyond network architectures, and the development of adaptive, AI-driven technologies for future intelligent communication systems.
Maaz Shafi received the B.Sc. degree in Electrical Engineering from the University of Engineering and Technology (UET), Mardan, Pakistan, in 2025. His academic interests include signals and systems, communication systems, digital signal processing, and advanced wireless and mobile communication systems. His recent research focuses on deep learning-based channel estimation techniques for 6G networks, where he modeled stochastic 6G wireless channels by incorporating reconfigurable intelligent surfaces (RIS) and massive MIMO techniques to generate data for supervised machine learning algorithms and successfully estimated channel statistics using deep neural networks. His other research interests include antenna design and communication, further exploration of artificial intelligence for wireless and mobile communication systems, and designing beamformers (i.e., precoders and combiners) for 6G and beyond networks while incorporating advanced technologies such as NOMA and NGMA for the optimization of future wireless communication systems.
Said Muhammad received his Bachelor’s degree in Electrical Engineering from the University of Engineering and Technology (UET), Mardan, Pakistan, in 2025. His academic interests primarily focus on advanced wireless communication technologies, with a special emphasis on machine learning and deep learning applications in next-generation communication systems. His recent research involves the development of deep learning-based channel estimation techniques for 6G networks. Said Muhammad is passionate about exploring innovative solutions in the field of artificial intelligence for wireless systems and aims to contribute to the advancement of intelligent, high-capacity, and energy-efficient communication technologies. His future goals include advancing research in MIMO, OFDM, and AI-powered wireless networks to shape the evolution of robust and intelligent communication systems for the 6G and beyond era.

### v2

@@ I Introduction
Reconfigurable intelligent surface (RIS) panels have emerged as a transformative technology for beyond-fifth-generation (B5G) and sixth-generation (6G) wireless systems. These panels are low-cost, nearly passive metasurfaces that can programmatically control the wireless propagation environment [1, 5]. RIS is especially beneficial in millimeter-wave (mmWave) and terahertz (THz) bands, where blockage and severe path loss are common. By intelligently redirecting reflections, RIS can extend coverage and improve spectral efficiency without requiring active radio-frequency (RF) chains on the surface [2, 3, 6]. RIS can also create a line-of-sight (LoS) link for users that are otherwise in non-LoS conditions. Such LoS links can further enhance localization accuracy by providing stronger and more stable reference signals [27, 28]. They also enable higher received signal strength and lower path loss, which improve link reliability and coverage at the cell edge [26]. In addition, robust LoS connectivity supports more flexible and dense deployments of RIS panels and access points in next-generation wireless systems. However, all these benefits critically rely on the acquisition of accurate channel state information (CSI) [4].
@@
A primary challenge is to obtain CSI for the cascaded link that involves the base station (BS), the RIS, and the user. Because RIS elements are passive, the BS–RIS and RIS–user channels cannot be estimated independently. Instead, only their cascaded product is observable at the receiver. In frequency-division duplexing (FDD) systems, where uplink-downlink reciprocity does not hold, downlink pilots must activate the RIS, and the user measures the combined reflected signals. For an RIS with $N$ elements and a BS with $M$ antennas, the cascaded channel comprises $NM$ complex parameters. Traditional least-squares (LS) estimation demands at least $Q\geq NM$ pilot symbols (or $Q\geq N^{\prime}M$ with grouping, where $N^{\prime}<N$), which becomes impractical for large $N$ due to excessive overhead [7, 8, 10, 11, 12]. Advanced estimators exploit channel structures, such as angular sparsity and spatial correlations, to reduce pilot requirements [4]. Deep neural networks can map low-dimensional pilot observations to the full high-dimensional cascaded channel, often integrating compressive sensing or denoising techniques [13, 14]. Nevertheless, models trained in isolated user or scenario-specific settings frequently exhibit poor generalization to varying user locations, blockage conditions, or angular spreads. This is exacerbated by non-independent and identically distributed (non-IID) data across a cell, leading to degraded accuracy and reliability [13, 14, 16]. Distributed machine learning (DML), including federated learning (FL), tackles non-IID data by aggregating model updates from distributed devices without centralizing raw data, as in the Federated Averaging (FedAvg) algorithm [19]. For heterogeneous environments, specialization enhances performance: sparsely-gated mixture-of-experts (MoE) models use a learned gate to route inputs to a subset of experts, balancing accuracy and efficiency [20]. In RIS CSI estimation, a gate based on coarse scenario indicators (e.g., elevation sectors) can select region-specific experts, avoiding generic feature extraction. Recent developments, such as transformer-based CSI prediction [17] and comprehensive surveys [16], emphasize the need for estimators that (i) function with few pilots ($Q$), (ii) generalize across regions, and (iii) maintain low online computational cost at user equipment (UE). In this paper, we present a DML framework for downlink cascaded channel estimation under limited pilot regimes. Our approach integrates three key components: (i) RIS grouping to compress the effective cascaded channel from $NM$ to $N^{\prime}M$ parameters (with $N^{\prime}=N/g$ for group size $g$), thereby reducing pilot and computational demands; (ii) a region-gated MoE architecture that classifies the propagation region first and then activates only the relevant expert for feature extraction from the pilot data; and (iii) synchronous FedAvg involving the BS and users to train a shared estimator on diverse, non-IID data without exchanging raw samples [19]. The gate employs sparse routing inspired by [20], ensuring that only one expert is active during inference, maintaining UE runtime comparable to a single compact convolutional neural network (CNN).
@@ I-A Related Work
To handle heterogeneity across users and regions without aggregating raw data, DML aggregates model parameters from devices using algorithms like FedAvg [19, 21]. In RIS contexts, [15] proposed a DML scheme for downlink channel estimation, demonstrating enhanced generalization over per-user training. Compared to [15], our work adds (i) RIS grouping to shrink the parameter space from $NM$ to $N^{\prime}M$, cutting pilot and compute costs; and (ii) a region-gated design that engages only one expert per inference, reducing on-device complexity while preserving accuracy.
@@ I-B Contributions
  • A synchronous FedAvg-based DML protocol involving users and the BS to train the shared model and experts on local non-IID data, enhancing generalization across regions [19, 15].
@@ II-D RIS Grouping Model
To reduce pilot overhead and computational complexity, we incorporate RIS element grouping via a fixed selection/averaging matrix $\mathbf{S}\in\{0,1\}^{N^{\prime}\times N}$ that maps $N$ physical elements to $N^{\prime}$ control units ($N^{\prime}=N/g$ for cluster size $g$). This grouping aggregates nearby elements into clusters, preserving total reflection energy while reducing the degrees of freedom [22]. The RIS phase vectors satisfy
$\boldsymbol{\phi}=\frac{1}{\sqrt{g}}\mathbf{S}^{\mathrm{T}}\bar{\boldsymbol{\phi}},\qquad\bar{\boldsymbol{\phi}}\in\mathbb{C}^{N^{\prime}\times 1},\ \ |\bar{\phi}_{i}|=1,$ (10)
@@
Throughout this work we model the RIS as having ideal continuous-valued phase control. Each element uses $\phi_{q,n}=e^{j\theta_{q,n}}$ with $\theta_{q,n}\in[0,2\pi)$ and $|\phi_{q,n}|=1$. This corresponds to an upper bound on the performance of practical RISs. Practical RISs often employ $B$-bit phase shifters with $\theta_{q,n}$ restricted to a finite set [23]. Explicit few-bit quantization constraints on $\{\theta_{q,n}\}$ into the pilot and beamforming design are left for future work. The linear observation model in (8) remains valid when $\phi_{q,n}$ is replaced by its quantized counterpart. The grouped variant in (12) also remains valid under quantization.
@@ II-E Problem Formulation
$\underbrace{\mathbf{y}_{r,k}^{\mathrm{p}}}_{\in\mathbb{C}^{Q}}=\underbrace{\boldsymbol{\Psi}_{r,k}}_{\in\mathbb{C}^{Q\times NM}}\underbrace{\mathbf{h}_{r,k}}_{\in\mathbb{C}^{NM}}+\ \mathbf{n}_{r,k}.$ (13)
@@
$\hat{\mathbf{h}}_{r,k}=g_{\boldsymbol{\vartheta}}\big(\mathbf{y}_{r,k}^{\mathrm{p}};\boldsymbol{\Psi}_{r,k}\big).$ (14)
@@
$\mathrm{NMSE}\triangleq\mathbb{E}\!\left[\frac{\|\hat{\mathbf{h}}_{r,k}-\mathbf{h}_{r,k}\|_{2}^{2}}{\|\mathbf{h}_{r,k}\|_{2}^{2}}\right].$ (15)
@@
$\hat{\mathbf{h}}_{r,k}^{\mathrm{LS}}=\big(\boldsymbol{\Psi}_{r,k}^{\mathrm{H}}\boldsymbol{\Psi}_{r,k}\big)^{-1}\boldsymbol{\Psi}_{r,k}^{\mathrm{H}}\mathbf{y}_{r,k}^{\mathrm{p}}.$ (18)
@@
MMSE requires a channel covariance model; we adopt a data-driven covariance estimated from samples in the baselines section ( Sec. IV).
@@ III-B3 Feature Mapper
$\hat{\mathbf{h}}_{r,k}=M_{\boldsymbol{\psi}}(\mathbf{Z}_{r,k}),$ (22)
@@ III-D Distributed Training Protocol
The DML training employs a synchronous federated learning protocol, with the BS serving as the central parameter server and the UEs as clients. This setup leverages the synchronous Federated Averaging (FedAvg) algorithm [19] to aggregate local updates, effectively handling the non-IID nature of channel data across regions. Let $\boldsymbol{\vartheta}=(\{\boldsymbol{\theta}_{r}\}_{r=1}^{R},\boldsymbol{\psi})$ denote the trainable parameters of the experts and mapper. In each communication round, the BS broadcasts the current model, clients perform local gradient computations on their private datasets, and the BS aggregates these updates using data-weighted averaging to update the global model.
@@
12:  Aggregate at BS with weights $p_{r,k}=|\mathcal{D}_{r,k}|/\sum_{r^{\prime},k^{\prime}}|\mathcal{D}_{r^{\prime},k^{\prime}}|$:
@@ III-F Computational Complexity
$\mathcal{C}_{\mathrm{map}}=\mathcal{O}\big(E_{1}E_{2}\cdot 32\cdot(2NM)\big),$ (27)
@@ A. Arrays, Pilots, and SNR
RIS pilots $\boldsymbol{\Theta}$ unit-modulus entries $\phi_{q,n}=e^{j\theta_{q,n}}$, $\theta_{q,n}\sim\mathrm{Unif}[0,2\pi)$22footnotemark: 2
@@ B. Channel Models ((2), (3))
  • Unit-modulus pilots satisfy $|\phi_{q,n}|=1$. For energy normalization, we use $\tilde{\mathbf{y}}^{\mathrm{p}}=\mathbf{y}^{\mathrm{p}}/\sqrt{Q}$ without altering (8).
@@ IV-C Ablations and Discussion
Distributed training (DML) vs. centralized/single-user training: synchronous gradient-averaging DML outperforms single-region or per-user training, with gains most evident at low SNR where regional differences in noise and angles are pronounced. This collaborative approach leverages non-IID data without privacy risks.
@@ Acknowledgment
The research team thanks the Deanship of Graduate Studies and Scientific Research at Najran University for supporting the research project through the Nama’a program, with the project code NU/GP/SERC/13/401-5.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Ethics Approval
This study did not involve human participants, animal subjects, or any identifiable personal data. All experiments were carried out using synthetic wireless channel models and Monte Carlo simulations of RIS-assisted 6G systems. No measurements were collected from operational networks, commercial devices, or real users.
Ethical approval was therefore not required, in line with the institutional guidelines for simulation-based research that does not involve human or animal subjects or personal data. No ethics committee or IRB reference number was issued for this work.
Author Contributions
Saifur Rahman: Conceptualization (equal); Data curation (equal); Software (equal); Validation (equal); Visualization (equal); Project administration (equal); Funding acquisition (equal). Syed Luqman Shah: Conceptualization (equal); Methodology (equal); Software (equal); Data curation (equal); Supervision (equal); Writing – original draft (equal); Validation (equal). Salman Khan: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Validation (equal); Writing – review & editing (equal). Jalal Khan: Data curation (equal); Formal analysis (equal); Investigation (equal); Validation (equal); Visualization (equal). Muhammad Irfan: Formal analysis (equal); Investigation (equal); Validation (equal); Project administration (equal); Funding acquisition (equal). Maaz Shafi: Formal analysis (equal); Investigation (equal); Validation (equal); Writing – review & editing (equal). Said Muhammad: Investigation (equal); Validation (equal); Visualization (equal); Writing – review & editing (equal). Fazal Muhammad: Data curation (equal); Investigation (equal); Methodology (equal); Supervision (equal); Visualization (equal); Writing – review & editing (equal). Mohammad Shahed Akond: Investigation (equal); Validation (equal); Visualization (equal); Project administration (equal); Funding acquisition (equal).
DATA AVAILABILITY
The data that support the findings of this study are available within the article.